Bioreactors are used in industry to semi-automate manufacturing processes involving microorganisms. In a small-scale pilot bioreactor using the plastic-degrading Pseudomonas strain, 10.0 L of 0.5% (m/v) 2,4-diaminotoluene (C7N2H10) reacts with at 8.0 L of O2 gas at 12.0 atm and 21.0oC.

A. To minimize contributing even more to climate change, the CO2 produced should ideally be stored and not released into the atmosphere. How many moles of CO2 gas will be produced by the small scale bioreactor?

B. If only 5.0 L of CO2 at 1 atm was actually captured, what would be the new pressure if the CO2 was compressed into at 1.0 L cylinder?

C. The ammonia produced could be used as a fertilizer for food production. Assuming the final volume is still 10 L, what is the concentration of ammonia produced by the small scale bioreactor?
1
Expert's answer
2020-12-22T04:30:59-0500




ANSWER1. Ag2​CrO4​⇌2Ag++CrO4​2−
Ksp=(2s)2×s=4s2
Ksp=(1.1×10−12)
S=34
Ksp​=6.5×10−5M
2. AgCl⇌Ag++Cl−
Ksp=S×S ;    Ksp=1.8×10−10
S=s=1.34×10−5M
3. AgBr⇌Ag++Br−
Ksp=S×S ;   Ksp=5×10−13
S=s=0.71×10−6M
4. AgI⇌Ag++I−
Ksp=S×S ;   Ksp=8.3×10−17
S=s=0.9×10−8M
∴ Solubility of Ag2​CrO4​ is highest, so it will precipitate last.